OpenStudy (pphalke):

the formula m=1/5 t can be used to determine how far M, in miles, you are from lightning when its thunder takes t seconds to reach your ears. if it takes 6 sec for the sound to reach you after you have seen the lightning, how far is the storm?

OpenStudy (anonymous):

m = 1/5t t = 6 so... 1/5(6) m = 6/5, or 1 1/5 so it is 1 1/5 miles away

OpenStudy (pphalke):

its 1 1/2 right?

OpenStudy (pphalke):

one and one fifth?

OpenStudy (anonymous):

1 and 1/5 miles x3